# Starburst-Driven Galactic Outflows

### v1

@@ 1 Introduction
In this study, we assess the role of an extended CR halo in modifying the development of galactic winds driven by CR and thermal gas pressure, and derive the criteria for the breakout of an outflow from a galaxy with a CR halo.
@@ 2 Starburst-driven Outflows in Galaxy Halos
The collective feedback from a central galactic starburst can initiate a blastwave, which may develop into a sustained galactic wind if the starburst activity is continuous. Typically, about one SN explodes per $\sim$100 M of star formation, with the exact rate dependent on the choice of the stellar initial mass function (e.g. Leitherer et al., 1999). We can therefore link the supernova event rate $\mathcal{R}_{\text{SN}}=\mathcal{R}_{-3}\,\text{kyr}^{-1}$ to the SFR of a galaxy by $\mathcal{R}_{\text{SF}}\sim 100\,\text{M}_{\odot}\,\mathcal{R}_{\text{SN}}$. Observationally, star formation activity is typically quantified using the SFR surface density, $\Sigma_{\text{SFR}}$. This can be related to $\mathcal{R}_{\text{SF}}$ by considering that most star formation contributing to the blastwave occurs within a cylindrical region with a radius comparable to the disk scale height, $\sim H_{\text{s}}$, of a galaxy (defined by eq. 8), i.e. $\Sigma_{\text{SFR}}=\mathcal{R}_{\text{SF}}\,/\left(\pi H_{\text{s}}^{2}\right)=\Sigma_{\text{SFR, 0}}\,\text{M}_{\odot}\,\text{yr}^{-1}\,\text{kpc}^{-2}$.
The injection rates of mass ($\dot{M}_{\text{SB}}=M_{\text{ej}}\mathcal{R}_{\text{SN}}$, for $M_{\text{ej}}=M_{\text{ej, 0}}\,\text{M}_{\odot}$ as the typical supernova ejecta mass) and energy ($\dot{E}_{\text{SB}}=E_{\text{SN}}\mathcal{R}_{\text{SN}}$, for $E_{\text{SN}}=10^{51}\,E_{51}\,\text{erg}$ as the typical mechanical energy supplied by a supernova) supplied to an expanding blastwave from a galactic starburst can be linked to the SFR through the mass- and energy-loading factors $\eta_{\text{m}}\equiv\dot{M}_{\text{SB}}/\mathcal{R}_{\text{SF}}=0.01M_{\text{ej, 0}}$ and $\eta_{\text{e}}\equiv\varepsilon_{\text{w}}\dot{E}_{\text{SB}}/\left(E_{\text{SN}}\mathcal{R}_{\text{SN}}\right)=\varepsilon_{\text{w}}$, respectively, where $\varepsilon_{w}$ is a thermalization efficiency factor accounting for energy dissipation in the system (see e.g. Thompson et al., 2016; Kim et al., 2017; Steinwandel et al., 2024). For convenience, we introduce the scaling parameters $\eta_{\text{m, -2}}=\eta_{\text{m}}/0.01$, $\eta_{\text{e, -2}}=\eta_{\text{e}}/0.01$ and $\varepsilon_{\text{w, -2}}=\varepsilon_{\text{w}}/0.01$. The supply of CRs to the wind is parametrized by $f_{\text{CR}}$, representing the CR energy fraction at the galactic mid-plane. In our model, these parameters are treated as constants, remaining fixed throughout the evolution of the outflow, and the flow is considered to be driven by the combination of central thermal and kinetic energy-injection, and CR pressure gradients.
@@ 2.1 CR halos and their effects on outflows
The formation of CR halos is a consequence of CR production during galaxy evolution. The long energy loss times of hadronic CRs in these environments (see Appendix A.1) ensure that most of the CR energy density supplied to a galaxy halo during its development can survive to the present day. Galaxies with significant historical stellar mass buildup are expected to host rich CR halos, even if their current star formation activity is low. Observations in $\gamma$-rays tentatively support this distinction, with CR halos primarily identified around massive late-type galaxies, while lower-mass galaxies show no indications of hosting such structures (Pshirkov & Nizamov, 2024).
@@
[…]
@@ 2.2 Outflow breakout criterion and terminal velocity
The terminal velocities predicted by our model allow for comparison with observations. Figure 2 shows terminal outflow velocities as a function of SFR surface density for a fiducial model with (dimensionless) parameters $n_{0}=1$, $\sigma_{1}=1$, $\eta_{\text{m, -2}}=5$, and varying values of $f_{\text{CR}}$ and $\eta_{\text{e, -2}}$ as indicated in the legend. These calculations assume that all galaxies in the sample share a similar dynamical equilibrium pressure, $P_{\text{DE}}\sim\rho_{\text{mp}}\sigma^{2}\sim G\Sigma^{2}$ (Ostriker & Kim, 2022) for $\rho_{\text{mp}}$ as the galactic mid-plane gas volume density and $\Sigma$ as the corresponding surface density that sets $\Sigma_{\text{SFR, c}}$. For comparison, observed outflow velocities for a sample of nearby starburst galaxies with stellar masses in the range $\log_{10}(M_{*}/\text{M}_{\odot})\in[7.1-10.9]$ are also shown (Heckman et al., 2015). Our model captures the general trend of observed flow velocities with parameter choices of $\varepsilon_{\text{w,-2}}\sim\eta_{\text{e,-2}}\sim 1$ and $\eta_{\text{m, -2}}\sim M_{\text{ej, 0}}\sim 5$. These choices align well with the energy and mass loading factors reported in numerical simulations of galactic outflows at altitudes of a few kpc (e.g. Kim & Ostriker, 2018; Rathjen et al., 2021; Steinwandel et al., 2024; Kjellgren et al., 2025).
@@ 3 Discussion and Implications
The structure of galactic winds has been extensively studied through theoretical approaches (e.g. Chevalier & Clegg, 1985; Fielding & Bryan, 2022; Modak et al., 2023), detailed numerical simulations (e.g. Kim & Ostriker, 2018; Vasiliev et al., 2023; Kjellgren et al., 2025) and observational studies (e.g. Krieger et al., 2019; Xu et al., 2023b; Bolatto et al., 2024). These studies have shown that galactic winds are ubiquitous, particularly among star-forming galaxies, and that galaxies with high SFRs tend to drive faster and hotter winds. However, not all galaxies show clear signatures of outflows. Some systems show that gas launched from the ISM is recycled within a galaxy ecosystem rather than expelled (Marasco et al., 2023). While the qualitative framework for wind launching is well established (see Thompson & Heckman, 2024, for a review), few studies quantify outflow launching conditions (e.g. Heckman et al., 2015; Orr et al., 2022).
Heckman et al. (2015) examined outflow speeds of galactic winds in a sample of nearby starbursts, and reported a sharp drop in flow velocities when the central SFR surface density fell below a critical threshold of $\Sigma_{\text{SFR, c}}\sim 1\,M_{\odot}\,\text{yr}^{-1}\,\text{kpc}^{-2}$. Our model shows that this critical SFR surface density arises due to the depth of the gravitational potential, which can only be overcome by sufficiently strong starbursts. The predicted value aligns with observations when considering typical dynamical equilibrium pressures set by the weight of the ISM, together with wind-loading parameters that are consistent with recent numerical studies of galactic winds in dwarf (Steinwandel et al., 2024) and spiral galaxies in the local Universe (Kjellgren et al., 2025). However, we note that other studies find significantly higher mass and energy loading, depending on the methodology used (e.g. Muratov et al., 2015; Smith et al., 2024).
@@
Most numerical simulations of starburst-driven galactic winds show that the inclusion of CRs improves their ability to drive warm outflows with high mass-loading factors (e.g. Girichidis et al., 2016; Rathjen et al., 2021; Chan et al., 2022; Armillotta et al., 2024). This is in agreement with our results, which indicate that outflow speed in the CR-dominated regime is independent of the mass-loading factor, and suggest that CR-driven winds can sustain substantial mass-loading before being significantly slowed. However, current numerical models do not typically include a pre-existing CR halo in their initial conditions. Instead, they only model the accumulation of CRs within a galaxy ecosystem over time, usually as a consequence of stellar feedback. The widespread findings of substantial feedback impacts from CR-driven outflows in the literature may reflect this limitation — a situation that would have certain parallels with earlier numerical studies that lacked CGM thermal pressure, leading to unrealistically strong outflows in simulations of isolated galaxies (Shin et al., 2021).
@@
At high redshift, several studies have reported high outflow velocities up to 1,000 $\text{km s}^{-1}$ (Sugahara et al., 2019; Xu et al., 2022). In general, higher outflow speeds require higher energy loading factors and lower mass loading factors. Observations suggest that high-redshift galaxies tend to be compact and turbulent (Genzel et al., 2023), with high gas fractions and surface densities (e.g. Genzel et al., 2011), as well as low metallicities (Maiolino et al., 2008). Since the critical SFR surface density scales with the dynamical equilibrium pressure as $\dot{\Sigma}_{\text{SFR, c}}\propto P_{\text{DE}}\propto\Sigma^{2}$, thermally-driven outflows become ineffective in a highly turbulent, high surface-density environment. On the other hand, lower metallicities result in longer cooling times, leading to higher momentum-loading (Oku et al., 2022). If mass-loading factors in high-redshift galaxies are comparable to those in low-redshift galaxies, then lower metallicities could explain the high observed outflow velocities in high-redshift galaxies, provided that the increased ISM weight does not suppress outflows.
@@
Curiously, the critical SFR surface density exhibits a stronger dependence on gas surface density than on the SFR surface density itself, which follows $\dot{\Sigma}_{\text{SFR}}\propto\Sigma^{1.4}$ (Kennicutt, 1989). This suggests that, counterintuitively, fast thermally-powered galactic outflows are expected to be more common at low surface densities, such as in dwarf galaxies, while being suppressed in extreme star-forming environments - including massive clumps in high-redshift galaxies (Genzel et al., 2011) and the proposed feedback-free starburst galaxies at $z\sim 10$ (Finkelstein et al., 2023; Dekel et al., 2023). At high redshifts, CR-driven slow outflows may be more prevalent in these systems, as there would not have been sufficient time for them to establish a CR halo capable of suppressing outflows. Indeed, highly mass-loaded outflows are essential for regulating star formation and explaining observed metallicity trends at high redshift (Toyouchi et al., 2025), which could naturally be accounted for by CR-driven outflows.
@@ 4 Conclusions
While our simplified approach is subject to substantial limitations (see Appendix B), it provides useful insights into the qualitative behavior of starburst-driven outflows and the influence of a CR halo. However, more detailed studies - including numerical simulations with CR halos as an initial condition - are needed to properly explore the physical impacts of CRs on the dynamical processes within galaxy halos.
@@ Acknowledgements.
This research was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. E.R.O is an international research fellow under the Postdoctoral Fellowship of the Japan Society for the Promotion of Science (JSPS), supported by JSPS KAKENHI Grant Number JP22F22327, and also acknowledges support from the RIKEN Special Postdoctoral Researcher Program for junior scientists. This work was supported in part by the MEXT/JSPS KAKENHI grant numbers 20H00180, 22K21349, 24H00002, and 24H00241 (K.N.). K.N. acknowledges the support from the Kavli IPMU, the World Premier Research Centre Initiative (WPI), UTIAS, the University of Tokyo.

### v2

@@ 1 Introduction
In this study, we assess the role of an extended CR halo in modifying the development of galactic winds driven by CR and thermal gas pressure, and derive the criteria for the breakout of an outflow from a galaxy with a CR halo. Our model focuses on the asymptotic, steady-state behavior of outflows driven by continuous central feedback, with particular emphasis on spatial scales extending from a few kiloparsecs above the galactic disk to $\sim$  100 kpc into the halo. The relevant timescales correspond to $\sim$ 10 Myr for shock breakout and $\sim 10-100$ Myr for the wind to approach its terminal velocity, depending on the dominant driving mechanism. While our approach does not capture the time-dependent, bursty nature of star formation or CR transport in evolving magnetic fields, it provides a physically motivated framework for understanding the long-term impact of CR halos on outflow dynamics.
@@ 2 Starburst-driven Outflows in Galaxy Halos
The collective feedback from a central galactic starburst can initiate a blastwave, which may develop into a sustained galactic wind if the starburst activity is continuous. Typically, about one SN explodes per $\sim$100 M of star formation, with the exact rate dependent on the choice of the stellar initial mass function (e.g. Leitherer et al., 1999). We can therefore link the supernova event rate $\mathcal{R}_{\text{SN}}=\mathcal{R}_{-3}\,\text{kyr}^{-1}$ to the SFR of a galaxy by $\mathcal{R}_{\text{SF}}\sim 100\,\text{M}_{\odot}\,\mathcal{R}_{\text{SN}}$. Observationally, star formation activity is typically quantified using the SFR surface density, $\Sigma_{\text{SFR}}$. This can be related to $\mathcal{R}_{\text{SF}}$ by considering that most star formation contributing to the blastwave occurs within a cylindrical region with a radius comparable to the disk scale height, $\sim H_{\text{s}}$, of a galaxy (defined by eq. 8), i.e. $\Sigma_{\text{SFR}}=\mathcal{R}_{\text{SF}}\,/\left(\pi H_{\text{s}}^{2}\right)=\Sigma_{\text{SFR, 0}}\,\text{M}_{\odot}\,\text{yr}^{-1}\,\text{kpc}^{-2}$. As we discuss in App. B, the time it takes to establish an outflow ranges from several 10s of Myr to several 100s of Myrs (see also e.g. Rathjen et al., 2023; Girichidis et al., 2024; Rathjen et al., 2025), while individual starburst episodes only last about 10 Myr (Rathjen et al., 2023). Thus it is important to note that the SFR entering our model should be interpreted as an average over many starburst episodes that collectively drive an outflow.
The injection rates of mass ($\dot{M}_{\text{SB}}=M_{\text{ej}}\mathcal{R}_{\text{SN}}$, for $M_{\text{ej}}=M_{\text{ej, 0}}\,\text{M}_{\odot}$ as the typical supernova ejecta mass) and energy ($\dot{E}_{\text{SB}}=E_{\text{SN}}\mathcal{R}_{\text{SN}}$, for $E_{\text{SN}}=10^{51}\,E_{51}\,\text{erg}$ as the typical mechanical energy supplied by a supernova) supplied to an expanding blastwave from a galactic starburst can be linked to the SFR through the mass- and energy-loading factors $\eta_{\text{m}}\equiv\dot{M}_{\text{SB}}/\mathcal{R}_{\text{SF}}=0.01M_{\text{ej, 0}}$ and $\eta_{\text{e}}\equiv\varepsilon_{\text{w}}\dot{E}_{\text{SB}}/\left(E_{\text{SN}}\mathcal{R}_{\text{SN}}\right)=\varepsilon_{\text{w}}$, respectively, where $\varepsilon_{w}$ is a thermalization efficiency factor accounting for energy dissipation in the system (see e.g. Thompson et al., 2016; Kim et al., 2017; Steinwandel et al., 2024). Similarly, we account for CR energy-dissipation with $\varepsilon_{\text{w, CR}}$ which for simplicity we here set to $\varepsilon_{\text{w, CR}}=\varepsilon_{\text{w}}=\eta_{\text{e}}$. For convenience, we introduce the scaling parameters $\eta_{\text{m, -2}}=\eta_{\text{m}}/0.01$, $\eta_{\text{e, -2}}=\eta_{\text{e}}/0.01$ and $\varepsilon_{\text{w, -2}}=\varepsilon_{\text{w}}/0.01$. The supply of CRs to the wind is parametrized by $f_{\text{CR}}$, representing the CR energy fraction at the galactic mid-plane. In our model, these parameters are treated as constants, remaining fixed throughout the evolution of the outflow, and the flow is considered to be driven by the combination of central thermal and kinetic energy-injection, and CR pressure gradients.
@@ 2.1 CR halos and their effects on outflows
The formation of CR halos is a consequence of CR production during galaxy evolution. The long energy loss times of hadronic CRs in these environments (see Appendix A.1) ensure that most of the CR energy density supplied to a galaxy halo during its development can survive to the present day. Galaxies with significant historical stellar mass buildup are expected to host rich CR halos, even if their current star formation activity is low. Observations in $\gamma$-rays tentatively support this distinction, with CR halos primarily reported around massive late-type galaxies, while lower-mass galaxies show no indications of hosting such structures (Pshirkov & Nizamov, 2024).
@@
where $\sigma=10\,\sigma_{1}\,\text{km/s}$ is the gas velocity dispersion, and $v_{\infty,2}$ is the rescaled terminal flow velocity defined as $v_{\infty}=100\,v_{\infty,2}\,\text{km/s}$ at large distances from the galactic plane. Equation 1 indicates that $z_{\text{CR}}$ is typically located at low altitudes for most galaxies. This suggests that CR-driven outflows are easily suppressed by the presence of an extended CR halo, if the halo has a scale-height $\gg H_{s}$. Galactic winds in systems with a well-developed CR halo are therefore expected to experience suppression, with their driving primarily dependent on thermal and kinetic energy-injection rather than CR pressure.
@@ 2.2 Outflow breakout criterion and terminal velocity
The terminal velocities predicted by our model allow for comparison with observations. Figure 2 shows terminal outflow velocities as a function of SFR surface density for a fiducial model with (dimensionless) parameters $n_{0}=1$, $\sigma_{1}=1$, $\eta_{\text{m, -2}}=5$, and varying values of $f_{\text{CR}}$ and $\eta_{\text{e, -2}}$ as indicated in the legend. These calculations assume that all galaxies in the sample share a similar dynamical equilibrium pressure, $P_{\text{DE}}\sim\rho_{\text{mp}}\sigma^{2}\sim G\Sigma^{2}$ (Ostriker & Kim, 2022) for $\rho_{\text{mp}}$ as the galactic mid-plane gas volume density and $\Sigma$ as the corresponding surface density that sets $\Sigma_{\text{SFR, c}}$. For comparison, observed outflow velocities for a sample of nearby starburst galaxies with stellar masses in the range $\log_{10}(M_{*}/\text{M}_{\odot})\in[7.1-10.9]$ are also shown (Heckman et al., 2015). Our model captures the general trend of observed flow velocities with parameter choices of $\varepsilon_{\text{w,-2}}\sim\eta_{\text{e,-2}}\sim 1$ and $\eta_{\text{m, -2}}\sim M_{\text{ej, 0}}\sim 5$. These choices align well with the energy and mass loading factors reported in numerical simulations of galactic outflows at altitudes of a few kpc (e.g. Kim & Ostriker, 2018; Rathjen et al., 2021; Steinwandel et al., 2024; Quataert & Hopkins, 2025).
@@ 3 Discussion and Implications
The structure of galactic winds has been extensively studied through theoretical approaches (e.g. Chevalier & Clegg, 1985; Fielding & Bryan, 2022; Modak et al., 2023), detailed numerical simulations (e.g. Kim & Ostriker, 2018; Vasiliev et al., 2023; Quataert & Hopkins, 2025) and observational studies (e.g. Krieger et al., 2019; Xu et al., 2023b; Bolatto et al., 2024). These studies have shown that galactic winds are ubiquitous, particularly among star-forming galaxies, and that galaxies with high SFRs tend to drive faster and hotter winds. However, not all galaxies show clear signatures of outflows. Some systems show that gas launched from the ISM is recycled within a galaxy ecosystem rather than expelled (Marasco et al., 2023). While the qualitative framework for wind launching is well established (see Thompson & Heckman, 2024, for a review), few studies quantify outflow launching conditions (e.g. Heckman et al., 2015; Orr et al., 2022).
Heckman et al. (2015) examined outflow speeds of galactic winds in a sample of nearby starbursts, and reported a sharp drop in flow velocities when the central SFR surface density fell below a critical threshold of $\Sigma_{\text{SFR, c}}\sim 1\,M_{\odot}\,\text{yr}^{-1}\,\text{kpc}^{-2}$. Our model shows that this critical SFR surface density arises due to the depth of the gravitational potential, which can only be overcome by sufficiently strong starbursts. The predicted value aligns with observations when considering typical dynamical equilibrium pressures set by the weight of the ISM, together with wind-loading parameters that are consistent with recent numerical studies of galactic winds in dwarf (Steinwandel et al., 2024) and spiral galaxies in the local Universe (Quataert & Hopkins, 2025). However, we note that other studies find significantly higher mass and energy loading, depending on the methodology used (e.g. Muratov et al., 2015; Smith et al., 2024).
@@
Most numerical simulations of starburst-driven galactic winds show that the inclusion of CRs improves their ability to drive warm outflows with high mass-loading factors (e.g. Girichidis et al., 2016; Rathjen et al., 2021; Chan et al., 2022; Armillotta et al., 2024). This is in agreement with our results, which indicate that outflow speed in the CR-dominated regime is independent of the mass-loading factor, and suggest that CR-driven winds can sustain substantial mass-loading before being significantly slowed. However, current numerical models do not typically include a pre-existing CR halo in their initial conditions. Instead, they only model the accumulation of CRs within a galaxy ecosystem over time, usually as a consequence of stellar feedback. The widespread findings of substantial feedback impacts from CR-driven outflows in the literature may reflect this limitation — a situation that has certain parallels with earlier numerical studies that lacked CGM thermal pressure, leading to unrealistically strong outflows in simulations of isolated galaxies (Shin et al., 2021).
We emphasize that CR-driven outflows are unlikely to operate efficiently in massive galaxies with deep gravitational potentials. Numerical studies (e.g. Jacob et al., 2018; Girichidis et al., 2024) show that in halos with $M_{\rm halo}\gtrsim 10^{12}\,M_{\odot}$, CR-driven winds are largely suppressed. This is due to a combination of strong gravitational binding, which raises the threshold for escape, and enhanced CR energy losses via Coulomb and hadronic interactions, which limit the dynamical role of CR pressure. These findings are consistent with our model, in which the outward CR pressure gradient becomes ineffective in the presence of an established CR halo, particularly when the halo’s scale height greatly exceeds the CR pressure equilibrium height. We thus interpret the CR suppression mechanism in our model as particularly relevant for lower-mass or early-phase starburst galaxies, while in more massive systems, thermal or radiation-driven mechanisms likely dominate the wind launching process.
@@
At high redshift, several studies have reported high outflow velocities up to 1,000 $\text{km s}^{-1}$ (Sugahara et al., 2019; Xu et al., 2022). In general, higher outflow speeds require higher energy loading factors and lower mass loading factors. Observations suggest that high-redshift galaxies tend to be compact and turbulent (Genzel et al., 2023), with high gas fractions and surface densities (e.g. Genzel et al., 2011), as well as low metallicities (Maiolino et al., 2008). Since the critical SFR surface density scales with the dynamical equilibrium pressure as $\dot{\Sigma}_{\text{SFR, c}}\propto P_{\text{DE}}\propto\Sigma^{2}$, thermally-driven outflows become ineffective in a highly turbulent, high surface-density environment. On the other hand, lower metallicities result in longer cooling times, leading to higher momentum loading (Oku et al., 2022). If mass-loading factors in high-redshift galaxies are comparable to those in low-redshift galaxies, then lower metallicities could explain the high observed outflow velocities in high-redshift galaxies, provided that the increased ISM weight does not suppress outflows.
@@
Curiously, the critical SFR surface density exhibits a stronger dependence on gas surface density than on the SFR surface density itself, which follows $\dot{\Sigma}_{\text{SFR}}\propto\Sigma^{1.4}$ (Kennicutt, 1989). This suggests that, counterintuitively, fast thermally-powered galactic outflows are expected to be more common at low surface densities, such as in dwarf galaxies, while being suppressed in extreme star-forming environments — including massive clumps in high-redshift galaxies (Genzel et al., 2011) and the proposed feedback-free starburst galaxies at $z\sim 10$ (Finkelstein et al., 2023; Dekel et al., 2023). At high redshifts, CR-driven slow outflows may be more prevalent in these systems, as there would not have been sufficient time for them to establish a CR halo capable of suppressing outflows. Indeed, highly mass-loaded outflows are essential for regulating star formation and explaining the observed metallicity trends at high redshift (Toyouchi et al., 2025), which can be naturally accounted for by CR-driven outflows.
We note that our model is intended as an idealized, steady-state framework that captures the asymptotic behavior of outflows under continuous feedback. While this does not resolve the detailed time-dependent evolution of bursty star formation or episodic feedback cycles, it provides a useful time-averaged description representative of either (i) long-duration starburst episodes or (ii) the cumulative effects of successive, shorter bursts over galactic dynamical timescales. A full treatment of the time dependence—including intermittent injection, CR streaming and diffusion, and adiabatic losses in 3D—will be essential in future work to extend the applicability of our model to more realistic galactic evolution scenarios.
@@ 4 Conclusions
While our simplified approach is subject to substantial limitations (see Appendix C), it provides useful insights into the qualitative behavior of starburst-driven outflows and the influence of a CR halo. However, more detailed studies - including numerical simulations with CR halos as an initial condition - are needed to properly explore the physical impacts of CRs on the dynamical processes within galaxy halos.
@@ Acknowledgements.
We thank the anonymous referee for their insightful comments and suggestions that helped to improve the quality of this work. This research was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. E.R.O is an international research fellow under the Postdoctoral Fellowship of the Japan Society for the Promotion of Science (JSPS), supported by JSPS KAKENHI Grant Number JP22F22327, and also acknowledges support from the RIKEN Special Postdoctoral Researcher Program for junior scientists. This work was supported in part by the MEXT/JSPS KAKENHI grant numbers 20H00180, 22K21349, 24H00002, and 24H00241 (K.N.). K.N. acknowledges the support from the Kavli IPMU, the World Premier Research Centre Initiative (WPI), UTIAS, the University of Tokyo.